OpenStudy (anonymous):

what is Solve for f. f + (–15) = –32 i need help plz

OpenStudy (shamil98):

f + (-15) = -32 f - 15 = -32 Think of it as this, take 15 away from f you get -32 to get f you want to undo that so you add 15 to both sides of this, what do you get? f - 15 + 15 = -32 + 15

OpenStudy (shamil98):

The idea here is to get f by itself.

OpenStudy (anonymous):

um ok thx

OpenStudy (anonymous):

thx

OpenStudy (shamil98):

Btw that needs to be simplifed

OpenStudy (shamil98):

anytime